a ball is kicked off the ground with a speed of 10m/s at an angel of 30* above horizontal. it lands on 3m net. find kinematic info

xi,yi

xf, yf

vix viy

vfx vif

ax ay

t
Answer #1

It lands on 3 m net ?????

is it the height at which the ball lands ????

If yes,

then

yf = 3 m

yi = 0 m ( launched from ground)

xi = 0 m

xf =

vix = 10 * cos 30 = 8.66 m/s

viy = 10 * sin 30 = 5 m/s

ax = 0 m/s2

ay = - 9.8 m/s2

vfx = 8.66 m/s (horizontal speed does not change at any time)

Now, to find time t,

yf = yi + vyt - 1/2gt2

3 = 0 + 5t - 4.9t2

we get a quadratic equation

Now, here comes the problem, as you can see this quadratic equation cannot be solved ....that's why please check you question again....is it 3 m final height or 3 m horizontal distance ??

4.9t2 - 5t + 3 = 0